40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 22 02:43:28 2011, kiwamu, Update, General, in-vac work 
    Reply  Mon Aug 22 21:32:10 2011, kiwamu, Update, General, more in-vac work : AS clipping fixed and OSEM/oplev adjustment 
       Reply  Tue Aug 23 06:49:24 2011, Anamaria, Update, General, more in-vac work : AS clipping fixed and OSEM/oplev adjustment 
          Reply  Tue Aug 23 09:40:37 2011, kiwamu, Update, General, Re: AS clipping fixed 
Message ID: 5275     Entry time: Mon Aug 22 02:43:28 2011     Reply to this: 5279
Author: kiwamu 
Type: Update 
Category: General 
Subject: in-vac work 

[Kieko / Kiwamu]

 The AS beam shows a little bit of clipping and right now this is the only concern for the alignment of the interferometer.

Other than that everything is okay including :

  + Beam centering on all the suspended optics

  + Arm cavity alignments. The Fabry-Perot fringes from both arm cavity were found on the AS camera.

  + POX/POY/POP, they are still successfully coming out from the chambers

  + IPPOS/ANG

  + Alignment of the green beams and the associated optics. Both green are reaching to the PSL table

 

We need to check/fix the AS beam clipping and once it's done we will readjust the OSEM mid range and the oplevs.

Then it will be ready for the drag/wipe and door closing.

ELOG V3.1.3-